# Optimization-Based Analysis of Music Intervals and Tuning Systems in Oral Traditions Using Pitch Histograms: A Case Study of Iranian Vocal Music

### v1

@@
Abstract
This paper presents a computational methodology for analyzing music intervals and tunings in microtonal oral traditions, utilizing pitch histograms, Dynamic Time Warping (DTW), and optimization techniques. By extracting pitch frequencies directly from vocal performances and aligning them with MIDI notes via DTW, we determine musical intervals using histograms. This approach offers an efficient, performance-based, and instrument-independent alternative to traditional tuning system analysis. Optimization techniques are then employed to align intervals throughout the oral tradition repertoire, capturing the specific tunings and modes involved. Our methodology demonstrates the potential of computational techniques in advancing musicological and ethnomusicological research, revealing new insights into the studied traditions.
Abstract
@@ 2 Choice of Repertoire
To demonstrate our methodology and examine its results, we have selected a comprehensive repertoire of 145 pieces from Iranian classical music, performed by one of the twentieth century’s masters, Karimi. These performances have been publicly recorded and transcribed by the prominent ethnomusicologist Masoudieh [1]. The transcriptions were later digitized in Finale, with pitch bends incorporated into the MIDI playback. This data has been compiled into a database of Iranian classical music as part of the first author’s dissertation [2].
Iranian vocalists are among the few musicians who continue to preserve and transmit music through oral tradition, relying on master-to-student transmission rather than musical notation. In the foreword to Marufi’s radif, Barkeshli emphasizes that the human voice provides a more authentic reference for Iranian musical intervals than musical instruments. Unlike instruments, which may introduce structural tuning constraints, the human voice remains unbounded, allowing for greater microtonal flexibility [3]. According to Barkeshli, measuring the Persian scale through vocal performances offers a more genuine representation of traditional intonation. Furthermore, in Iranian classical music, the human voice serves as the primary pitch reference, with instruments—especially unfretted ones like the kamancheh and ud—seeking to closely follow vocal intonation patterns.
@@
Figure 1: Transcription of Example 1
@@ 3 Pitch Recognition
To generate the pitch file, we used Sonic Annotator with the following parameters: a step size of 256, a block size of 2048, low amplitude suppression of 0.1, onset sensitivity of 0.7, prune threshold of 0.1, and a threshold distribution of 2. Pitch values extracted in Hertz (Hz) were converted to the cents system, where each octave equals 1200 cents. The interval between two notes with frequencies $f_{1}$ and $f_{2}$ in cents is given by the following formula.
$1200\times\log_{2}(\frac{f_{2}}{f_{1}})$ (1)
@@ 4.1 Pitch Histogram
Histograms have been applied extensively to non-Western and microtonal musical traditions, providing insights into unique interval patterns and tuning systems. Serra et al. used histograms to analyze the maqam system in Arab music, identifying specific microtonal intervals that define this tradition [6]. Similarly, Srinivasamurthy et al. utilized histograms to study raga structures in Indian classical music, revealing distinct interval patterns that differentiate ragas from each other [7]. Barış Bozkurt has significantly contributed to the use of pitch histograms in the analysis of Turkish makam music [8]. Bozkurt’s work emphasizes the importance of computational methods for analyzing the microtonal intervals characteristic of makam music, providing a deeper understanding of its intricate tuning systems.
@@
Figure 2 presents the pitch histogram of the audio recording of Example 1. To identify the peaks, we first determine the range of each mountain. This process is challenging, as peak boundaries are not always clearly defined. We employ different methods based on both the shape of the mountain and the derivative of the smoothed histogram. Tracking sign changes in the derivative curve provides an indication of peak boundaries. Figure 3 illustrates this process for determining the range of the mountain corresponding to the note F in Example 1.
Figure 2: Pitch Histogram of Example 1
Figure 3: Determining the Range of note F in Example 1
@@ II. Hidden Secondary Peaks
Figure 4: Types of Pitch Histogram Peaks
@@ 5 Alignment of Pitch and MIDI
Figure 5: Aligned Pitch and MIDI
@@ 6 Histogram of Notes
Figure 6: Note Histogram of Example 1
@@ 7 Computational Analysis of Intonation across the Repertoire
After identifying all the peaks in the pitch histograms, we compile them into a table, where each row represents the peaks corresponding to a specific piece. For the selected repertoire, we visualize the first category of pieces in Figure 7, while the full dataset consists of 145 gushes. The complete tables are available in the corresponding GitHub repository for reference.
Figure 7: All the Observed Peaks for the first 14 Pieces of the Repertore
In the selected repertoire, the reference pitch is fluid, meaning that there is no fixed fundamental pitch across all pieces, and the intervals between notes can vary depending on the performance. For example, the interval between F and G might be 198 cents in one piece but 210 cents in another. This variability makes it challenging to align multiple pieces. To address this, we first approximate the alignment of all pieces, forming columns of corresponding notes. However, their placement will not initially be optimal. We iteratively refine their alignment by shifting the notes of each piece up or down while minimizing the total sum of standard deviations across corresponding notes, that is explained in Section 7.1.
@@ 7.1 Optimization Process for Aligning Pitch Rows in the Tuning System
Figure 8: Evolution of the Cost Function During the Optimization Process
@@ 7.2 Derived Tuning and Visualization of Intonation Patterns
Figure 9 shows the final tuning extracted from the multiple pieces in shur dastgah.
Figure 9: Tuning Extracted from Dastgāh of shur
8 acknowledgments

### v2

# Computational Extraction of Intonation and Tuning Systems from Multiple Microtonal Monophonic Vocal Recordings with Diverse Modes

@@
Computational Extraction of Intonation and Tuning Systems from Multiple Microtonal Monophonic Vocal Recordings with Diverse Modes
Sepideh Shafiei Shapour Hakam missing
Аннотация
@@
Author’s version (preprint).
This paper appears in the Proceedings of the Sound and Music Computing Conference (SMC 2025).
© 2025 The Authors. Licensed under CC BY 4.0.
Please cite the published version: https://zenodo.org/records/15838081
@@ 2 Choice of Repertoire
To demonstrate our methodology and examine its results, we have selected a comprehensive repertoire of 145 pieces from Iranian classical music, performed by one of the twentieth century’s masters, Karimi. These performances have been publicly recorded and transcribed by the prominent ethnomusicologist Masʿudiye [1]. The transcriptions were later digitized in Finale, with pitch bends incorporated into the MIDI playback. This data has been compiled into a database of Iranian classical music as part of the first author’s dissertation [2].
Iranian vocalists are among the few musicians who continue to preserve and transmit music through oral tradition, relying on master-to-student transmission rather than musical notation. In the foreword to Maʿrufi’s radif, Barkeshli emphasizes that the human voice provides a more authentic reference for Iranian musical intervals than musical instruments. Unlike instruments, which may introduce structural tuning constraints, the human voice remains unbounded, allowing for greater microtonal flexibility [3]. According to Barkeshli, measuring the Persian scale through vocal performances offers a more genuine representation of traditional intonation. Furthermore, in Iranian classical music, the human voice serves as the primary pitch reference, with instruments—especially unfretted ones like the kamancheh and ʿud—seeking to closely follow vocal intonation patterns.
@@
In standard MIDI convention, each pitch is represented by an integer; for example, 55 corresponds to G2, 56 to G#2, and 57 to A2. To accommodate microtonal intervals such as sori (half-sharp) and koron (half-flat), we extend this system by doubling the MIDI resolution. Specifically, we multiply each MIDI pitch number by two, enabling us to insert intermediary microtonal steps. For example, in our system, 110 represents G2, 111 corresponds to G2-half-sharp (sori), 112 to G#2, 113 to A2-half-flat (koron), and 114 to A2.
Рис. 1: Transcription of Example 1
@@ 3 Pitch Recognition
To generate the pitch file, we used Sonic Annotator with the following parameters: a step size of 256, a block size of 2048, low amplitude suppression of 0.1, onset sensitivity of 0.7, prune threshold of 0.1, and a threshold distribution of 2. All audio files were sampled at 44.1 kHz. Pitch values extracted in Hertz (Hz) were converted to the cents system, where each octave equals 1200 cents. The interval between two notes with frequencies $f_{1}$ and $f_{2}$ in cents is given by the following formula:
$\text{cents}(f_{1},f_{2})=1200\cdot\log_{2}\left(\frac{f_{2}}{f_{1}}\right)$
While the fixed block size provides a consistent basis for our analyses, we acknowledge that certain aspects of microtonal vocal performance—particularly rapid expressive gestures—might benefit from multiresolution or adaptive time-frequency representations. Exploring such approaches remains a promising direction for future development.
@@ 4.1 Pitch Histogram
Histograms have been applied extensively to non-Western and microtonal musical traditions, providing insights into unique interval patterns and tuning systems. Serra used histograms to analyze the maqam system in Arab music, identifying specific microtonal intervals that define this tradition [6]. Similarly, Srinivasamurthy et al. utilized histograms to study raga structures in Indian classical music, revealing distinct interval patterns that differentiate ragas from each other [7]. Barış Bozkurt has significantly contributed to the use of pitch histograms in the analysis of Turkish makam music [8]. Bozkurt’s work emphasizes the importance of computational methods for analyzing the microtonal intervals characteristic of makam music, providing a deeper understanding of its intricate tuning systems.
@@
Figure 2 presents the pitch histogram of the audio recording of Example 1. To identify the peaks, we first determine the range of each mountain. This process is challenging, as peak boundaries are not always clearly defined. We employ different methods based on both the shape of the mountain and the derivative of the smoothed histogram. Tracking sign changes in the derivative curve provides an indication of peak boundaries. Figure 3 illustrates this process for determining the range of the mountain corresponding to the note F in Example 1. In Figures 2 and 3, we only annotate peaks that are supported by a sufficient number of data points to ensure reliable estimation. Although a small peak corresponding to G is present, its low height suggests limited data coverage, making accurate identification and interpretation less robust. For this reason, such minor peaks are not labeled or further analyzed in this study.
Рис. 2: Pitch Histogram of Example 1
Рис. 3: Determining the Range of note F in Example 1
@@ II. Hidden Secondary Peaks
Рис. 4: Types of Pitch Histogram Peaks
@@ 5 Alignment of Pitch and MIDI
Рис. 5: Aligned Pitch and MIDI
@@ 6 Histogram of Notes
Рис. 6: Note Histogram of Example 1
@@ 7 Computational Analysis of Intonation across the Repertoire
After identifying all the peaks in the pitch histograms, we compile them into a table, where each row represents the peaks corresponding to a specific piece. For the selected repertoire, we visualize the first category of pieces (shur) in Figure 7, while the full dataset consists of 145 gushes. The complete tables are available in the corresponding GitHub repository for reference.
Рис. 7: All the Observed Peaks for the first 14 Pieces of the Repertoire
In the selected repertoire, the reference pitch is fluid, meaning that there is no fixed fundamental pitch across all pieces, and the intervals between notes can vary depending on the performance. For example, the interval between F and G might be 198 cents in one piece but 210 cents in another. This variability makes it challenging to align multiple pieces.
To further illustrate the diversity of tuning behaviors in the dataset, we examined several examples across different dastgāhs. In particular, we compared the intervals derived from audio performances with those found in musicological transcriptions of the same pieces. We observed notable inconsistencies, especially in the treatment of neutral intervals. For example, what is labeled as a "neutral second"in some cases may span anywhere from approximately 120 to 180 cents. In certain contexts, an interval of 180 cents has been labeled a neutral second, whereas in others it has been considered a major second. Similarly, intervals around 120 cents have been interpreted either as narrow neutral seconds or as minor seconds, depending on the transcriber’s theoretical framework. These discrepancies highlight both the flexibility of intonation in performance and the lack of consensus in interval classification, reinforcing the typology described in Section 4.2 and the need for data-driven representations.
To address this, we first approximate the alignment of all pieces, forming columns of corresponding notes. However, their placement will not initially be optimal. We iteratively refine their alignment by shifting the notes of each piece up or down while minimizing the total sum of standard deviations across corresponding notes, that is explained in Section 7.1.
@@ 7.1 Optimization Process for Aligning Pitch Rows in the Tuning System
The current optimization procedure is based on a greedy iterative search that ensures monotonic descent of the cost function. This approach offers interpretability and low computational overhead, but we acknowledge that its local, discrete adjustments may limit the exploration of the solution space and lead to premature convergence. As such, this method is viewed as a baseline. In future work, we plan to explore gradient-based and probabilistic optimization strategies, which may offer more robust convergence properties and improve tuning alignment, particularly in cases involving sparse or unevenly distributed data.
Рис. 8: Evolution of the Cost Function During the Optimization Process
@@ 7.2 Derived Tuning and Visualization of Intonation Patterns
Figure 9 shows the final tuning extracted from the multiple pieces in shur. It is not intended as a prescriptive or canonical tuning system for this dastgāh, but rather as a descriptive statistical summary of how pitches are realized across multiple pieces. While some intervals may align with theoretical expectations or commonly used tunings, others reflect the fluidity and expressive variation inherent in this repertoire. Therefore, this tuning should be understood as a performance-informed approximation, not a fixed template.
Рис. 9: Tuning Extracted from Dastgāh of shur
8 Future Directions
The methodology introduced in this study opens several promising avenues for future work. One natural extension is to apply the analysis across a broader range of vocalists within the same genre, in order to uncover stylistic characteristics of intonation that may distinguish individual performers, lineages, or pedagogical traditions. By systematically comparing pitch distributions and tuning profiles across singers, it may be possible to identify performer-specific tendencies or regionally influenced variants of modal interpretation.
In addition to stylistic analysis, this approach can also be used to explore the nuanced realizations of ornamentation and intervallic ambiguity that are central to microtonal vocal traditions. For instance, by examining how certain intervals (such as neutral seconds) are treated across different contexts, we may gain insight into their expressive functions and contextual elasticity.
Acknowledgments
@@
Список литературы
  • [1] M. T. Masudiye. The Vocal Radif of Traditional Iranian Music According to the version of Mahmud Karimi. Tehran: Māhur. 2004.
  • [2] S. Shafiei, Extracting Theory from Practice: A Computational Analysis of the Persian Radif. PhD dissertation at the Graduate Center, City University of New York. 2021.
  • [3] M. Maʿrufi. 2010. Radif-e haft dastgāh-e musiqi-e Irani.Tehran: Māhur.
  • [4] E. Gómez, M. Blaauw, J. Bonada, P. Chandna, and H. Cuesta. “Deep Learning for Singing Processing: Achievements, Challenges and Impact on Singers and Listeners."arXiv preprint. arXiv:1807.03046. 2018.
  • [5] B. Gfeller, C. Frank, D. Roblek, M. Sharifi, M. Tagliasacchi, and M. Velimirović, "SPICE: Self-Supervised Pitch Estimation."IEEE/ACM Transactions on Audio, Speech, and Language Processing 28 (2020): 1118-1128.
  • [6] X. Serra, 2017. The computational study of a musical culture through its digital traces. Acta Musicologica, 89(1), pp.24-44.
  • [7] A. Srinivasamurthy, and X. Serra, 2014, May. A supervised approach to hierarchical metrical cycle tracking from audio music recordings. In 2014 IEEE International Conference on Acoustics, Speech and Signal Processing (ICASSP) (pp. 5217-5221). IEEE.
  • [8] B. Bozkurt, 2008. "An automatic pitch analysis method for Turkish maqam music."Journal of New Music Research.
  • [9] B. Bozkurt. 2011. “Pitch Histogram Based Analysis of Makam Music in Turkey.” Paper presented at Les Corpus de l’oralité Colloquium, Strasbourg.
  • [10] J. Salamon, and E. Gomez, 2012. "Melody extraction from polyphonic music signals using pitch contour characteristics."IEEE Transactions on Audio, Speech, and Language Processing.
  • [11] G. Peeters, 2006. "Musical key estimation using spectral analysis techniques."Proceedings of the International Society for Music Information Retrieval Conference.
  • [12] T. Berg-Kirkpatrick, J. Andreas, and D. Klein, 2014. "Unsupervised transcription of piano music."Proceedings of the Conference on Empirical Methods in Natural Language Processing.
  • [13] S. Shafiei, S. Hakam, and A. Nick. 2024. "Quantifying Pitch Drift In Unaccompanied Solo Singing: A Computational Examination Through Density-Based Clustering."In Sound Music Conference Proceedings.
  • [14] S. Shafiei, 2019. "Analysis of Vocal Ornamentation in Iranian Classical Music."In Sound Music Conference Proceedings.
  • [15] S. Ewert, M. Muller, V. Konz, D. Mullensiefen, and G.A. Wiggins. "Towards Cross-Version Harmonic Analysis of Music."2012. IEEE Transactions on Multimedia 14(3): 770-782.
  • [16] M. Muller, S. Ewert, 2004. "Towards timbre-invariant audio features for harmony-based music."Proceedings of the IEEE International Conference on Acoustics, Speech, and Signal Processing.
  • [17] C. Raffel, and D. PW Ellis. "Optimizing DTW-based Audio-to-MIDI Alignment and Matching."In IEEE International Conference on Acoustics, Speech and Signal Processing (ICASSP): 81-85. 2016.